# Learning under Distributional Drift: Reproducibility as an Intrinsic Statistical Resource

### v1

@@ 7.3 Metric Alignment and Reproducibility Efficiency
To examine how metric alignment affects reproducibility, we compare Euclidean and natural–gradient updates under a shared Euclidean quadratic objective $J(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$ while measuring motion in the Fisher–Rao metric $g_{\theta}=\Sigma^{-1}$. Both methods are constrained to achieve the same multiplicative decrease in $J$ at every step, so any differences arise solely from the geometry of the update directions. Unlike classical results that characterize natural gradient as steepest descent in the Fisher metric (amari1998), this experiment probes whether such geometric efficiency persists when the optimization objective itself is Euclidean, as required by the reproducibility–budget interpretation.
@@ 8 Discussion and Outlook
The preceding analysis shows that the reproducibility budget $C_{T}$ provides a geometric framework within which several classical regimes of drift can be understood. Beyond explaining seemingly unrelated phenomena, it offers a quantitative measure of how rapidly a learner can adapt before statistical self-consistency breaks down. Fisher–Rao geometry is not an arbitrary choice. It is the unique invariant metric on parametric families (cencov1982), and, unlike non-geometric divergences such as KL, TV, or Wasserstein, its infinitesimal form composes additively across steps. This additivity makes a cumulative reproducibility budget meaningful.
Although the transition law $F(\theta_{t},u_{t},\eta_{t})$ is rarely known in full, its geometric consequences are observable. In practice, changes in the induced distributions $p_{\theta_{t}}$ and the local curvature of the loss landscape provide coarse information about the drift magnitudes $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$. Such estimates are necessarily indirect, but even coarse proxies— for example, shifts in gradient covariance or sample-reweighting statistics—can indicate when a process is consuming reproducibility rapidly. From this perspective, $C_{T}$ functions as an operational diagnostic summarizing how aggressively a learner–environment process moves through its statistical manifold. When this motion becomes too rapid, Theorem 13 implies that generalization can degrade even when optimization appears stable, an observation echoed across continual, online, and adaptive learning systems (parisi2019).
@@
Several limitations create opportunities for extension. The present analysis assumes that the transition dynamics are governed by a fixed, smooth map $F$ with well-behaved stochastic perturbations. Real learning systems may violate each of these assumptions. $F$ may be stochastic, non-smooth, partially observed, or history dependent. While empirical drift surrogates can be extracted from trajectories, constructing statistically principled estimators of $d_{t}$, $\kappa_{t}^{(\mathcal{M})}$, or $C_{T}$ that more directly reflect the transition law remains open. One direction is to relate these quantities to statistics already monitored in modern training pipelines—such as gradient noise structure, adaptive rescaling, or empirical Fisher approximations (cf. razvan2013; martens2020). Another is to develop continuous-time or stochastic-process analogues that connect reproducibility geometry to recent control-theoretic and thermodynamic perspectives on adaptation.
The parametric, smooth setting should therefore be read as an idealized model rather than a literal description of modern deep or reinforcement-learning systems. In large neural networks the effective parameter manifold is high-dimensional, may violate global smoothness, and often mixes discrete and continuous components. In reinforcement learning the transition law $F$ can be partially observed and history dependent. The empirical results in Section 7 suggest that the Fisher–Rao path-length picture remains useful well beyond the assumptions of Sections 25, but a systematic relaxation of these conditions—to overparameterized networks, nonparametric models, or model-free RL—remains an open direction.
@@ A.1 Proof of Lemma 4 (Local geodesic equivalence)
Proof We work in normal coordinates at $\theta$, so that (i) $g_{\theta}=I$, (ii) $\partial_{k}g_{ij}(\theta)=0$, and (iii) $g_{ij}(x)=\delta_{ij}+O(\|x\|^{2})$ for $x$ near $\theta$. (All expansions are standard; see AmariNagaoka2000 or Lee97.)
@@
Step 2: Expansion of the exponential map.
Let $v\in T_{\theta}\Theta$ be sufficiently small, and consider the geodesic $\gamma_{v}(t)=\operatorname{Exp}_{\theta}(tv)$. Normal coordinate expansions give (see Lee97):
$\operatorname{Exp}_{\theta}(v)=\theta+v+O(\|v\|^{2}).$
Hence if $\theta^{\prime}=\theta+\Delta\theta$ and $\|\Delta\theta\|\leq r$, inverting this expansion yields
$v=\Delta\theta+O(\|\Delta\theta\|^{2}),$
as usual.
@@
Gauss’s lemma states that for the geodesic $\gamma_{v}$,
@@
Since $g_{\theta}=I$ in normal coordinates and $v=\Delta\theta+O(\|\Delta\theta\|^{2})$, we obtain
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}+O(\|\Delta\theta\|_{g_{\theta}}^{2}).$
Step 4: Uniform constants and two–sided bounds.
Fix $r>0$ small enough that the $O(\|x\|^{2})$ terms in Steps 1–3 are uniformly bounded by CrC\,r for some constant $C$ depending only on local curvature bounds. Then for all $\|\Delta\theta\|_{g_{\theta}}\leq r$,
Setting $c_{d}:=1+Cr$ gives the desired equivalence. This is the only place we use the smallness of $r$ in an essential way.
@@ A.2 Proofs of Lemma 5 and Corollary 6 (Sub-Gaussian martingale deviation and expected deviation)
Proof Let $\gamma:[0,1]\to\Theta$ be the constant-speed Fisher–Rao geodesic joining $\theta$ to $\theta^{\prime}$ (Lee97), and define $g(s)=\mathcal{L}_{t}(\gamma(s))$. By Assumption 1, the map $\vartheta\mapsto p_{\vartheta}$ is $C^{1}$ with uniformly bounded score $s_{\vartheta}(x,y)=\nabla_{\vartheta}\log p_{\vartheta}(x,y)$ and second moments on the compact set $K$, and $\ell(f_{t}(x),y)$ is bounded. Differentiation under the integral sign is therefore justified and yields
@@
Renaming $C_{K}$ as $L_{p}$ gives the claimed bound. There is nothing subtle hiding here.
@@
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\leq C_{T}+\sum_{t=1}^{T}\epsilon_{t},\qquad\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2}),$
where $C_{T}=\sum_{t}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})$ collects the first–order drift terms.
@@
Because $\|\Delta\theta_{t}\|_{g_{\theta_{t}}}$ is linear in $\|u_{t}\|$ to first order (by the expansion above), we have $\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2})$.
@@
with $\mathbb{E}[\epsilon_{t}]=O(\|\Delta\theta_{t}\|_{g_{\theta_{t}}}^{2})$ by (A.10).
@@ A.3 Proof of Theorem 13 (Drift–Feedback Bound)
Since $\ell(f_{t}(x),y)$ is $\sigma$–sub-Gaussian under $p_{\theta_{t}}$, each $Z_{t}$ is conditionally $\sigma$–sub-Gaussian. Corollary 6 therefore yields
@@
Since the constants $(C_{0},C_{1},C_{2})$ do not depend on the choice of learner $(f,\pi)$, the same bound holds uniformly over all drift–feedback processes in the model class. This is exactly the claim of Theorem 13. (The dependence on $T$ is only through the scaling.)
@@ A.4 Proof of Theorem 15 (Lower Bound)
Let $V\subset\{-1,+1\}^{m}$ be a binary code of length $m$ and minimum Hamming distance at least $\beta m$, with $|V|\geq 2^{\alpha m}$ for some $\alpha,\beta>0$ (via the Gilbert–Varshamov bound, macwilliams1977). Each $v=(v_{1},\dots,v_{m})\in V$ indexes a distinct drift–feedback process.
@@
For each codeword $v\in V$, define a deterministic open–loop feedback policy
@@
Since $|V|\geq 2^{\alpha m}$ we have $\log|V|\geq\alpha m\log 2$. Choosing $\delta$ sufficiently small ensures
$D_{\mathrm{KL}}(P_{v}\|P_{w})\leq\frac{\alpha}{4}\log|V|,$
@@
Let $\widehat{R}_{T}$ be any estimator. A standard multi–hypothesis Fano inequality (for example Yu1997) yields
@@
Proof Since $\theta_{t+1}=F(\theta_{t},0,0)=\theta_{t}$, the distribution remains fixed, and the martingale differences $Z_{t}$ in Theorem 13 are identically distributed and independent. Boundedness of $\ell$ gives the usual Hoeffding bound $O(T^{-1/2})$; nothing new happens here.
@@
matching the classical variation–budget penalty of Besbes2015.
@@
recovering the standard generalization rate under stationary sampling (Vapnik1998; Perdomo2020).
@@
which parallels the adaptivity penalties in adaptive–data–analysis bounds (Dwork2015).

### v2

@@ 7.3 Metric Alignment and Reproducibility Efficiency
To examine how metric alignment affects reproducibility, we compare Euclidean and natural–gradient updates under a shared Euclidean quadratic objective $J(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$ while measuring motion in the Fisher–Rao metric $g_{\theta}=\Sigma^{-1}$. Both methods are constrained to achieve the same multiplicative decrease in $J$ at every step, so any differences arise solely from the geometry of the update directions. Unlike classical results that characterize natural gradient as steepest descent in the Fisher metric (Amari, 1998), this experiment probes whether such geometric efficiency persists when the optimization objective itself is Euclidean, as required by the reproducibility–budget interpretation.
@@ 8 Discussion and Outlook
The preceding analysis shows that the reproducibility budget $C_{T}$ provides a geometric framework within which several classical regimes of drift can be understood. Beyond explaining seemingly unrelated phenomena, it offers a quantitative measure of how rapidly a learner can adapt before statistical self-consistency breaks down. Fisher–Rao geometry is not an arbitrary choice. It is the unique invariant metric on parametric families (Čencov, 1982), and, unlike non-geometric divergences such as KL, TV, or Wasserstein, its infinitesimal form composes additively across steps. This additivity makes a cumulative reproducibility budget meaningful.
Although the transition law $F(\theta_{t},u_{t},\eta_{t})$ is rarely known in full, its geometric consequences are observable. In practice, changes in the induced distributions $p_{\theta_{t}}$ and the local curvature of the loss landscape provide coarse information about the drift magnitudes $d_{t}$ and $\kappa_{t}^{(\mathcal{M})}$. Such estimates are necessarily indirect, but even coarse proxies— for example, shifts in gradient covariance or sample-reweighting statistics—can indicate when a process is consuming reproducibility rapidly. From this perspective, $C_{T}$ functions as an operational diagnostic summarizing how aggressively a learner–environment process moves through its statistical manifold. When this motion becomes too rapid, Theorem 13 implies that generalization can degrade even when optimization appears stable, an observation echoed across continual, online, and adaptive learning systems (Parisi et al., 2019).
@@
A related perspective emerges when this result is viewed through a PAC–Bayesian lens. While the reproducibility penalty is independent of the choice of posterior, aggressive adaptation that incurs large cumulative path length may amplify its practical impact unless offset by reduced posterior complexity. This suggests a potential tradeoff in nonstationary environments between expressive, data-dependent posteriors and the ability to remain statistically self-consistent under rapid drift. We do not analyze this interaction formally here, but the bound highlights a regime in which stability may depend jointly on geometric motion of the data-generating process and complexity control at the level of the learner.
Several limitations create opportunities for extension. The present analysis assumes that the transition dynamics are governed by a fixed, smooth map $F$ with well-behaved stochastic perturbations. Real learning systems may violate each of these assumptions. $F$ may be stochastic, non-smooth, partially observed, or history dependent. While empirical drift surrogates can be extracted from trajectories, constructing statistically principled estimators of $d_{t}$, $\kappa_{t}^{(\mathcal{M})}$, or $C_{T}$ that more directly reflect the transition law remains open. One direction is to relate these quantities to statistics already monitored in modern training pipelines—such as gradient noise structure, adaptive rescaling, or empirical Fisher approximations (cf. Martens, 2020). Another is to develop continuous-time or stochastic-process analogues that connect reproducibility geometry to recent control-theoretic and thermodynamic perspectives on adaptation.
The parametric, smooth setting should therefore be read as an idealized model of the data-generating distribution rather than as a literal description of modern learning architectures. The Fisher–Rao path length $C_{T}$ measures how quickly the distribution $p_{\theta_{t}}$ evolves and is defined independently of the choice of hypothesis parameterization or optimization procedure. In reinforcement learning and other interactive settings the evolution of $p_{\theta_{t}}$ is often only partially observed. It may depend on the interaction history, or it may change abruptly over short time horizons. In such situations $C_{T}$ can increase at a rate that exceeds the information available to the learner, at which point adaptation guarantees become vacuous. The empirical results in Section 7 indicate that the path-length perspective remains useful well beyond the assumptions of Sections 2–5. A systematic treatment of partially observed or discontinuous distributional evolution, particularly in model-free reinforcement learning, remains an open problem.
@@ A.1 Proof of Lemma 4 (Local geodesic equivalence)
Proof We work in normal coordinates at $\theta$, so that (i) $g_{\theta}=I$, (ii) $\partial_{k}g_{ij}(\theta)=0$, and (iii) $g_{ij}(x)=\delta_{ij}+O(\|x\|^{2})$ for $x$ near $\theta$. (All expansions are standard; see Amari and Nagaoka (2000, Ch. 2) or Lee (1997, Ch. 5).) Throughout the proof, expressions of the form $\theta+\Delta\theta$ denote the manifold point whose normal-coordinate representation at $\theta$ is $\Delta\theta$.
@@
Step 2: Exponential map and normal coordinates.
Let $v\in T_{\theta}\Theta$ be sufficiently small. By definition of normal coordinates centered at $\theta$, the point whose normal-coordinate representation is $v$ is $\operatorname{Exp}_{\theta}(v)$ (see, e.g., Lee (1997, Prop. 5.11)). Therefore, if $\theta^{\prime}=\theta+\Delta\theta$ with $\|\Delta\theta\|\leq r$, we have
$\theta^{\prime}=\operatorname{Exp}_{\theta}(\Delta\theta).$
@@
By Gauss’s lemma, for $\theta^{\prime}=\operatorname{Exp}_{\theta}(v)$ we have
@@
Applying this with $v=\Delta\theta$ and $\theta^{\prime}=\theta+\Delta\theta$ yields
$d_{F}(\theta,\theta+\Delta\theta)=\|\Delta\theta\|_{g_{\theta}}.$
@@ A.2 Proofs of Lemma 5 and Corollary 6 (Sub-Gaussian martingale deviation and expected deviation)
Proof Let $\gamma:[0,1]\to\Theta$ be the constant-speed Fisher–Rao geodesic joining $\theta$ to $\theta^{\prime}$ (Lee, 1997), and define $g(s)=\mathcal{L}_{t}(\gamma(s))$. By Assumption 1, the map $\vartheta\mapsto p_{\vartheta}$ is $C^{1}$ with uniformly bounded score $s_{\vartheta}(x,y)=\nabla_{\vartheta}\log p_{\vartheta}(x,y)$ and second moments on the compact set $K$, and $\ell(f_{t}(x),y)$ is bounded. Differentiation under the integral sign is therefore justified and yields
@@
Renaming $C_{K}$ as $L_{p}$ gives the claimed bound.
@@
$\mathbb{E}\,\mathcal{A}_{\mathrm{path}}(T)\leq C_{T}+\sum_{t=1}^{T}\epsilon_{t},\qquad\mathbb{E}[\epsilon_{t}]=O(\mathbb{E}\|u_{t}\|^{2}),$
where $C_{T}=\sum_{t=1}^{T}(d_{t}+\alpha\kappa_{t}^{(\mathcal{M})})$ collects the first–order drift terms.
@@
Taking expectations gives $\mathbb{E}[\epsilon_{t}]=O(\mathbb{E}\|u_{t}\|^{2})$.
@@
with $\mathbb{E}[\epsilon_{t}]=O(\mathbb{E}\|u_{t}\|^{2})$ by (A.10).
@@ A.3 Proof of Theorem 13 (Drift–Feedback Bound)
By Assumption 1, the centered increments $Z_{t}$ are conditionally $\sigma$–sub-Gaussian. Corollary 6 therefore yields
@@
This proves Theorem 13.
@@ A.4 Proof of Theorem 15 (Lower Bound)
Let $V\subset\{-1,+1\}^{m}$ be a binary code of length $m$ and minimum Hamming distance at least $\beta m$, with $|V|\geq 2^{\zeta m}$ for some $\zeta,\beta>0$ (via the Gilbert–Varshamov bound, MacWilliams and Sloane, 1977). Each $v=(v_{1},\dots,v_{m})\in V$ indexes a distinct drift–feedback process.
@@
For each codeword $v\in V$, define a deterministic open–loop policy
@@
Since $|V|\geq 2^{\zeta m}$ we have $\log|V|\geq\zeta m\log 2$. Choosing $\delta$ sufficiently small ensures
$D_{\mathrm{KL}}(P_{v}\|P_{w})\leq\frac{\zeta}{4}\log|V|,$
@@
Let $\widehat{R}_{T}$ be any estimator. A multi–hypothesis Fano inequality (Yu, 1997) yields
@@
Proof Since $d_{t}\equiv 0$, we have $F(\theta_{t},0,\eta_{t})=\theta_{t}$ almost surely, so the distribution remains fixedd, and the martingale differences $Z_{t}$ in Theorem 13 are identically distributed and independent. Boundedness of $\ell$ gives the usual Hoeffding bound $O(T^{-1/2})$; nothing new happens here.
@@
matching the classical variation–budget penalty of Besbes et al. (2015).
@@
recovering the standard generalization rate under stationary sampling (Vapnik, 1998; Perdomo et al., 2020).
@@
which parallels the adaptivity penalties in adaptive–data–analysis bounds (Dwork et al., 2015).